OpenStudy (zela101):

How to graph this bounded region in 3d.

OpenStudy (zela101):

G is the solid in the first octant that is bounded by the parabolic cylinder z=2-x^2 and the plane z=0, y=x, and y=0.

OpenStudy (zela101):

This question consist of an integral to solve, but that's not the problem. I want to set up the limits, but to do that i have to draw so i can get an idea what intervals are present.

OpenStudy (mathmale):

"First octant" implies that x, y and z are all zero or greater.

OpenStudy (zela101):

Correct.

OpenStudy (mathmale):

"bounded by the parabolic cylinder z=2-x^2" Note that y is not mentioned here, meaning that y can have any value, but z (the vertical component of any point on the region) is a function of x with z-intercept (0,0,2). Looking at your imaginary solid along the y-axis, you should see an upside down parabola. Again, y can have any value.

OpenStudy (zela101):



OpenStudy (zela101):